Working on Algebra homework, I can figure out how to start this problem, but by the time I get to the end, my answer never makes sense.
2/y+3 - 1/y-3 = 1
I found the LCD, (y+3)(y-3), and multiplied that by each term. Then I cancelled out whatever I could and set up an equation:
2(y-3) - (y+3) = (y+3)(y-3)
Then I simplified that, but I ended up with y=y2
Thanks.

Answer by scott8148(6628)

0 = y^2 - y

factoring ___ 0 = y(y - 1)

0 = y

0 = y - 1 ___ 1 = y

if you do a thought experiment, one and zero are the only numbers which equal their own square (or any power)